# Learning Vision-Based Neural Network Controllers with Semi-Probabilistic Safety Guarantees

### v1

@@ Abstract
Ensuring safety in autonomous systems with vision-based control remains a critical challenge due to the high dimensionality of image inputs and the fact that the relationship between true system state and its visual manifestation is unknown. Existing methods for learning-based control in such settings typically lack formal safety guarantees. To address this challenge, we introduce a novel semi-probabilistic verification framework that integrates reachability analysis with conditional generative adversarial networks and distribution-free tail bounds to enable efficient and scalable verification of vision-based neural network controllers. Next, we develop a gradient-based training approach that employs a novel safety loss function, safety-aware data-sampling strategy to efficiently select and store critical training examples, and curriculum learning, to efficiently synthesize safe controllers in the semi-probabilistic framework. Empirical evaluations in X-Plane 11 airplane landing simulation, CARLA-simulated autonomous lane following, and F1Tenth lane following in a physical visually-rich miniature environment demonstrate the effectiveness of our method in achieving formal safety guarantees while maintaining strong nominal performance. Our code is available at https://github.com/xhOwenMa/SPVT.
I Introduction
@@
Many real-world applications, such as self-driving cars and robotic navigation, require controllers that process high-dimensional image inputs to make real-time decisions. The centrality of visual inputs (particularly when other modalities are limited or unreliable) thus makes ensuring the safety of vision-based control an important problem in trustworthy AI. However, verifying the safety of such controllers remains a major open challenge due to the complexity of image-based inputs and the high computational cost of traditional verification methods [1, 2, 3].
@@
While reinforcement learning (RL) with high-dimensional image inputs has shown promise in learning control policies that optimize performance [4, 5], most methods lack formal guarantees of safety [6]. Moreover, verifying the safety of neural network controllers operating in high-dimensional observation spaces remains computationally intractable. Existing approaches to safe control primarily focus on low-dimensional state inputs [7] and empirical safety evaluations. Other approaches, such as verified safety over the entire input region and control barrier function-based methods, have also been explored [8, 9]. However, these methods struggle when the controller operates on image inputs due to the high dimensionality of the observation space. Moreover, while dynamic behavior of many autonomous systems of interest has established models approximating their trajectories through the system state space, physics that map state to its visual representation are considerably more complex, and associated models far more involved and less reliable.
@@
In this work, we integrate reachability analysis with generative modeling to enable efficient verification of neural network controllers operating on high-dimensional image spaces [10, 11]. Specifically, we employ a conditional generative adversarial network (cGAN) [12] to model the perceptual mapping from states to images, allowing us to verify safety properties in a structured and lower-dimensional latent space. To address the scalability challenges of verification, we introduce a semi-probabilistic verification (SPV) framework, where safety properties are verified over a sampled distribution of initial states (using distribution-free tail bounds) but for all possible latent environment representations of the cGAN. In addition, we present a training algorithm that makes use of a novel safety loss as a differentiable proxy to this verification objective. A key component of this algorithm is our approach to adapt the training set, stochastically biasing it towards states for which safety is difficult to verify. As our experiments in simulated plane landing, as well as both simulated and physical autonomous lane following, demonstrate, the proposed approach yields control policies that exhibit considerably stronger safety properties compared to state-of-the-art safe vision-based control baselines.
@@
  • A novel semi-probabilistic safety verification framework (SPV) approach that provides formal safety guarantees while remaining computationally feasible in high-dimensional vision-based control settings.
@@
  • Experimental evaluation in simulated and physical lane following that demonstrates effective empirical and provably-verified performance of the policies trained through our approach in comparison with several state-of-the-art baselines.
II Related Work
Formally verifying the safety of vision-based controllers is very challenging. Traditional verification tools [2, 3, 13, 14] are too computationally demanding to scale to realistic networks. Recent progress have enabled verification of larger neural networks through over-approximation [15, 16, 17], and abstract interpretation [18, 19, 20]. However, verifying vision-based controllers remains a difficult task since it is inherently less straightforward to define safety properties in the image space.
A practical approach to verify vision-based controllers is by approximating the perception module. This reduces verification complexities by projecting the problem from image space to lower-dimensional state space [10, 11, 21]. These methods, however, only focus on post-hoc verification and cannot be integrated into the training process for learning a safer controller. For a more comprehensive survey on the verification of vision-based controllers, see [22].
Safe reinforcement learning encodes safety based on the constrained Markov decision processes [23] but it is often too soft to enforce strict safety constraints [24]. On the other hand, control-theoretic methods such as barrier functions [7, 9, 25], and reachability analysis [26, 27] are more powerful but they have limited scalability to high-dimensional system such as vision-based controllers. Recently, some works extend these methods to the image space [28, 29, 30, 31]. However, they function more like a safety check around the controller, giving yes/no answers for passed in reference control, without providing formal safety guarantees about the controller itself.
III Model
III-A Problem Formulation
@@
Let $P$ denote a safety specification, which is a predicate $P(s)$ indicating whether a state $s\in\mathcal{S}$ is safe or not. Similarly, let $P(\mathcal{T})$ indicate whether $P(s)$ is true for all $s\in\mathcal{T}\subseteq\mathcal{S}$. We assume that both $P(s)$ and $P(\mathcal{T})$ can be evaluated efficiently (for example, safety is often described using linear inequalities and $\mathcal{T}$ is a polyhedron). Given a policy $\pi$, the controlled dynamical system effectively becomes $s_{t+1}=f(s_{t},\pi(h(s_{t}))$. We say that this dynamical system is safe for an initial state $s_{0}$ over a horizon $K$ if $P(s_{t})$ is true for all $0\leq t\leq K$.
This notion of safety, however, is limited for two reasons. First, we do not know $h$, so we cannot directly verify the dynamical system above. Second, we wish for a policy $\pi$ to satisfy safety in a way that is not tied to a specific starting state, but with respect to the full set of initial states $\mathcal{S}$. We address the first challenge by leveraging a conditional generative adversarial network (cGAN) to approximate $h$, and the second by using a semi-probabilistic verification (SPV) framework. We describe both of these ideas next.
III-B Approximating the Visual Observation Model
We address the first challenge by using a conditional generative model $g(s,z)$ which induces a distribution over observations $o\in\mathcal{O}$ for a given state $s\in\mathcal{S}$, with $z\in\mathcal{Z}$ a (typically uniformly distributed) random vector, analogous to the approach proposed by Katz et al. [10]. Such a generator can be trained, for example, using the conditional generative adversarial network (cGAN) framework [12, 32] from a collection of data $(o,s)$ in which images $o$ are annotated with associated states $s$. We can view the latent random vectors $z$ as representations of natural environment variation (e.g., different perspectives, lighting, etc). The goal here is that $g$ approximates $h$, but in practice this assumption is too strong. Instead, we make the following considerably weaker assumption about the relationship between $h$ and $g$.
Assumption 1
@@ III-B Approximating the Visual Observation Model
In practice, this assumption boils down to having (a) sufficient training data for the generator $g$ and (b) a sufficiently rich representation (e.g., neural network) and latent dimension of $z$.
IV Semi-Probabilistic Verification
Our notion of safety is based on $K$-reachability. The principal distinction is that in the latter, safety is guaranteed for all states in some specified set $\mathcal{S}_{0}\subseteq\mathcal{S}$ from which dynamics may be initialized. However, the resulting $K$-reachability proofs are generally conservative and typically suffer from significant scalability challenges. When controllers use vision, scalability can be a prohibitive barrier to verification. In practice, however, we can often obtain information about the distribution over initial states of the dynamical system $s_{0}$. For example, by collecting empirical visual data and annotating it with state-relevant information (for example, Waymo [33] or KITTI [34] datasets in the case of autonomous driving). On the other hand, the distribution over the initial state is often difficult to cleanly characterize (indeed, it may be heavy-tailed). It is, therefore, natural to appeal to distribution-free bounds to obtain probabilistic safety proofs with respect to the unknown distribution $\mathcal{D}$ over $s_{0}$ based on safety properties obtained for a finite sample of initial states. In contrast, the distribution of the visual environment induced by $\omega$ is far more challenging to characterize or sample, particularly since we do not know $h$.
@@ IV Semi-Probabilistic Verification
To formalize, fix a policy $\pi$ and let $\mathcal{S}_{t+1}(s_{0},\pi)=\{f(s,\pi(o))|o=h(s,\omega),s\in\mathcal{S}_{t}(s_{0},\pi),\omega\in\Omega\},$ where $\mathcal{S}_{0}(s_{0},\pi)=\{s_{0}\}$. Define $\mathrm{Reach}_{K}(s_{0},\pi)=\cup_{t=0}^{k}\mathcal{S}_{t}(s_{0},\pi)$, that is, all states that can be reached from $s_{0}$ for any perceptual environment $\omega\in\Omega$. Note that this form of reachability cannot be verified, since we do not know $h$. However, we can now leverage the cGAN $g$ as a proxy, with Assumption 1 allowing us to obtain sound safety guarantees. Specifically, let
$\hat{\mathcal{S}}_{t+1}(s_{0},\pi)=\{f(s,\pi(o))|$ $o\in g(s,z)\pm\epsilon,$
@@
Proof:
Suppose $\mathcal{S}_{t}\subseteq\hat{\mathcal{S}}_{t}$ and let $s\in\mathcal{S}_{t}$ and $o=h(s,\omega)$ for some $\omega\in\Omega$. Then $s\in\hat{\mathcal{S}}_{t}(s_{0},\pi)$ and by Assumption (1), $o\in g(s,z)$ for some $z\in\mathcal{Z}$. Consequently, $f(s,\pi(o))\in\hat{\mathcal{S}}_{t+1}(s_{0},\pi)$. Since $\hat{\mathcal{S}}_{0}(s_{0},\pi)=\mathcal{S}_{0}(s_{0},\pi)=\{s_{0}\}$, the result follows by induction. ∎
@@
Theorem IV.2
Suppose that $\{s_{i}\}_{i=1}^{N}$ i.i.d. from $\mathcal{D}$ and let $V=\{s_{i}|P(\mathrm{Reach}_{K}(s_{i},\pi,g))\}$. Then under Assumption (1),
@@
with probability at least $1-\delta$
Proof:
Let $\alpha=\Pr_{s\sim\mathcal{D}}\left[P(\mathrm{Reach}_{K}(s,\pi,g))\right]$ and $\hat{\alpha}=\frac{|V|}{N}$. By the Chernoff-Hoeffding bound, $\Pr\left(|\hat{\alpha}-\alpha|\geq\epsilon\right)\leq 2e^{-2N\epsilon^{2}}$, where the probability is with respect to datasets of $N$ initial states. Letting $\delta=2e^{-2N\epsilon^{2}}$, we obtain the confidence bound: $\Pr\left(\alpha\geq\hat{\alpha}-\sqrt{\frac{1}{2N}\log\frac{2}{\delta}}\right)\geq 1-\delta$. Finally, since by Theorem IV.1, $P(\mathrm{Reach}_{K}(s_{0},\pi,g))\Rightarrow P(\mathrm{Reach}_{K}(s_{0},\pi))$, the result follows. ∎
@@
V Learning-Based Synthesis of Provably Safe Vision-Based Control
@@
The previous section shows that this still enables rigorous semi-probabilistic verification. Additionally, we consider a special case in which safety properties are tied to a scalar safety score (for example, cross-track error in lane following). In particular, let $\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$ be a safety score function over the reachable set, with $P(\mathcal{T})$ translating the safety score into a predicate (e.g., error exceeds a predefined threshold). We assume that we can obtain differentiable bounds on $\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$ (e.g., if we use $\alpha,\beta$-CROWN [15, 17, 35]).
V-A The Learning Framework
@@
A central challenge in synthesizing a provably safe policy $\pi_{\theta}$ in our setting arises from the involvement of high-dimensional images generated by the generator $g$ (as a proxy for the true perception model $h$), which serve as inputs to the controller. Our overall approach is as follows. First, we begin with a controller $\pi_{\hat{\theta}}$ that is empirically safe (e.g., pre-trained with a safe RL method), which we also use as an anchor controller to avoid sacrificing too much empirical performance as we train for safety verification. Next, starting with $\hat{\theta}$, we train (or fine-tune) $\pi_{\theta}$ to minimize $\sum_{i\in S_{l}}\mathcal{L}(s_{i},\theta),$ where $S_{l}$ is a set of initial states $s_{0}$ used in training which evolves over training iterations $l$ and $\mathcal{L}(\theta)$ an appropriate loss function. The central algorithm design questions thus amount to 1) the choice of the loss function, and 2) the problem of selecting data $S_{l}$ to use for training in each iteration, so as to ultimately obtain a provably (rather than merely empirically) safe policy. We address these questions next.
V-B Loss Function
@@
We propose a loss function that integrates both a supervised learning loss and safety loss as follows:
$\mathcal{L}(s_{i},\theta)=\lambda_{1}\mathcal{L}_{acc}(s_{i},\theta)+\lambda_{2}\mathcal{L}_{safety}(s_{i},\theta).$ (4)
@@
The supervised loss $\mathcal{L}_{acc}$ is the mean squared error loss between predictions of our controller and the anchor controller, and aims to preserve a strong empirical performance with respect to the original pre-trained anchor controller.
@@
Turning next to the safety loss, recall that we assume that safety is quantified by a safety score function $\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$. One candidate would simply be to use this score as part of the loss function. However, this is impractical, as it is typically intractable to compute at scale and to the extent that it can be done, the tools for doing so are not differentiable. However, neural network verification techniques exist which compute differentiable sound upper and lower bounds on this quantity, and these therefore make natural candidates to use in constructing a loss function. More precisely, let $\underline{\sigma}(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))\leq\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))\leq\overline{\sigma}(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$ (i.e., $\underline{\sigma}(\cdot)$ is the lower and $\overline{\sigma}(\cdot)$ the upper bound on $\sigma(\cdot)$). To simplify notation, we let $\sigma^{(i)}_{K}=\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$, with $\overline{\sigma}^{(i)}_{K}$ and $\underline{\sigma}^{(i)}_{K}$ the corresponding upper and lower bounds. Then we can define the safety loss as
@@
V-C Adaptive Training Data
Our adaptive training procedure performs gradient updates by sampling batches from an adaptive collection of training data $S_{l}$ which consists of two disjoint and fixed-size components: the set of random initial states $S_{0}$, and $S_{A}$, maintained as a priority queue, containing initial states for which safety is a challenge to satisfy. Specifically, when training starts, $S_{A}$ is empty, and we gradually populate $S_{A}$ during the warmup period by adding the $m$most challenging datapoints (in the sense detailed below) from each training batch to it. To ensure $S_{0}$ and $S_{A}$ remain disjoint, whenever a datapoint is added to $S_{A}$, it is also deleted from $S_{0}$, and we generate another datapoint uniformly randomly to add to $S_{0}$.

### v2

@@ Abstract
Ensuring safety in autonomous systems with vision-based control remains a critical challenge due to the high dimensionality of image inputs and the fact that the relationship between true system state and its visual manifestation is unknown. Existing methods for learning-based control in such settings typically lack formal safety guarantees. To address this challenge, we introduce a novel semi-probabilistic verification framework that integrates reachability analysis with conditional generative networks and distribution-free tail bounds to enable efficient and scalable verification of vision-based neural network controllers. Next, we develop a gradient-based training approach that employs a novel safety loss function, safety-aware data-sampling strategy to efficiently select and store critical training examples, and curriculum learning, to efficiently synthesize safe controllers in the semi-probabilistic framework. Empirical evaluations in X-Plane 11 airplane landing simulation, CARLA-simulated autonomous lane following, F1Tenth vehicle lane following in a physical visually-rich miniature environment, and Airsim-simulated drone navigation and obstacle avoidance demonstrate the effectiveness of our method in achieving formal safety guarantees while maintaining strong nominal performance.
Codehttps://github.com/xhOwenMa/SPVT
1 Introduction
@@
Many real-world applications, such as self-driving cars and robotic navigation, require controllers that process high-dimensional image inputs to make real-time decisions. The centrality of visual inputs (particularly when other modalities are limited or unreliable) thus makes ensuring the safety of vision-based control an important problem in trustworthy AI. However, verifying the safety of such controllers remains a major open challenge due to the complexity of image-based inputs and the high computational cost of traditional verification methods (huang2019reachnn; katz2017reluplex; huang2017safety).
@@
While reinforcement learning (RL) with high-dimensional image inputs has shown promise in learning control policies that optimize performance (le2022deep; kiran2021deep), most methods lack formal guarantees of safety (kochdumper2023provably). Moreover, verifying the safety of neural network controllers operating in high-dimensional observation spaces remains computationally intractable. Existing approaches to safe control primarily focus on low-dimensional state inputs (dawson2022safe) and empirical safety evaluations. Other approaches, such as verified safety over the entire input region and control barrier function-based methods, have also been explored (wu2024verified; emam2022safe). However, these methods struggle when the controller operates on image inputs due to the high dimensionality of the observation space. Moreover, while dynamic behavior of many autonomous systems of interest has established models approximating their trajectories through the system state space, physics that map state to its visual representation are considerably more complex, and associated models far more complex and less reliable.
@@
In this work, we integrate reachability analysis with generative modeling to enable efficient verification of neural network controllers operating on high-dimensional image spaces (katz2022verification; cai2024scalable). Specifically, we employ a generative neural network with latent inputs representing environmental and perceptual variation to model the perceptual mapping from states to images, allowing us to verify safety properties in a structured and lower-dimensional latent space. To address the scalability challenges of verification, we introduce a semi-probabilistic verification (SPV) framework, where safety properties are verified over a sampled distribution of initial states (using distribution-free tail bounds) but for all possible latent environment representations of the trained generative perception model. In addition, we present a training algorithm that makes use of a novel safety loss as a differentiable proxy to this verification objective. A key component of this algorithm is our approach to adapt the training set, stochastically biasing it towards states for which safety is difficult to verify. Our experiments in simulated plane landing, simulated and physical autonomous lane following, and drone navigation and obstacle avoidance, demonstrate that the proposed approach yields policies that exhibit significantly stronger safety properties compared to state-of-the-art safe control baselines.
@@
  • A novel semi-probabilistic safety verification (SPV) framework that provides formal safety guarantees while remaining computationally feasible in high-dimensional vision-based control settings.
@@
  • Experimental evaluation in simulated and physical path following and simulated drone control that demonstrates effective empirical and provably-verified performance of the policies trained through our approach in comparison with state-of-the-art baselines.
2 Model
2.1 Problem Formulation
@@
Let $P$ denote a safety specification, which is a predicate $P(s)$ indicating whether a state $s\in\mathcal{S}$ is safe or not. Similarly, let $P(\mathcal{T})$ indicate whether $P(s)$ is true for all $s\in\mathcal{T}\subseteq\mathcal{S}$. We assume that both $P(s)$ and $P(\mathcal{T})$ can be evaluated efficiently (for example, safety is often described using linear inequalities and $\mathcal{T}$ is a polyhedron). Given a policy $\pi$, the controlled dynamical system effectively becomes $s_{t+1}=f(s_{t},\pi(h(s_{t},\omega))$. We say that this dynamical system is safe for an initial state $s_{0}$ over a horizon $K$ if $P(s_{t})$ is true for all $0\leq t\leq K$.
This notion of safety, however, is limited for two reasons. First, we do not know $h$, so we cannot directly verify the dynamical system above. Second, we wish for a policy $\pi$ to satisfy safety in a way that is not tied to a specific starting state, but with respect to the full set of initial states $\mathcal{S}$. We address the first challenge by leveraging a conditional generative neural network model to approximate $h$, and the second by using a semi-probabilistic verification (SPV) framework. We describe both of these ideas next.
2.2 Approximating the Visual Observation Model
We address the first challenge by using a conditional generative model $g(s,z)$ which induces a distribution over observations $o\in\mathcal{O}$ for a given state $s\in\mathcal{S}$, with $z\in\mathcal{Z}$ a (typically distributed according to a uniform or Gaussian distribution) random vector, analogous to the approach proposed by katz2022verification. Such a generator can be trained, for example, using the conditional generative adversarial network (cGAN) framework (mirza2014conditional; isola2017image) or a conditional diffusion model (yang2023lossy) from a collection of data $(o,s)$ in which images $o$ are annotated with associated states $s$. We can view the latent random vectors $z$ as representations of natural environment variation (e.g., different perspectives, lighting, etc). The goal here is that $g$ approximates $h$, but in practice this assumption is too strong. Instead, we make the following considerably weaker assumption about the relationship between $h$ and $g$.
Assumption 1.
@@ 2.2 Approximating the Visual Observation Model
In practice, this assumption boils down to having (a) sufficient training data for the generator $g$ and (b) a sufficiently rich representation (e.g., neural network) and latent dimension of $z$. We validate this assumption in Section 5 below.
3 Semi-Probabilistic Verification
Our notion of safety is based on $K$-reachability. In traditional $K$-reachability, safety is guaranteed for all states in some specified set $\mathcal{S}_{0}\subseteq\mathcal{S}$ from which dynamics may be initialized. However, such $K$-reachability proofs are generally conservative and typically suffer from significant scalability challenges. When controllers use vision, scalability can be a prohibitive barrier to verification. In practice, however, we can often obtain information about the distribution $\mathcal{D}$ over initial states $s_{0}$ of the dynamical system. For example, by collecting empirical visual data and annotating it with state-relevant information (Waymo (sun2020scalability) or KITTI (geiger2013vision) datasets in the case of autonomous driving). On the other hand, the distribution over the initial state is often difficult to cleanly characterize (indeed, it may be heavy-tailed). It is, therefore, natural to appeal to distribution-free bounds to obtain probabilistic safety proofs with respect to the unknown distribution $\mathcal{D}$ over $s_{0}$ based on safety properties obtained for a finite sample of initial states. In contrast, the distribution of the visual environment induced by $\omega$ is far more challenging to characterize or sample, particularly since we do not know $h$.
@@ 3 Semi-Probabilistic Verification
To formalize, fix a policy $\pi$ and let
$\mathcal{S}_{t+1}(s_{0},\pi)=\{f(s,\pi(o))|$ $o=h(s,\omega),$
$s\in\mathcal{S}_{t}(s_{0},\pi),\omega\in\Omega\},$
where $\mathcal{S}_{0}(s_{0},\pi)=\{s_{0}\}$. Define $\mathrm{Reach}_{K}(s_{0},\pi)=\cup_{t=0}^{K}\mathcal{S}_{t}(s_{0},\pi)$, that is, all states that can be reached from $s_{0}$ for any perceptual environment $\omega\in\Omega$. Note that this form of reachability cannot be verified, since we do not know $h$. However, we can now leverage the generator $g$ as a proxy, with Assumption 1 allowing us to obtain sound safety guarantees. Specifically, let
$\hat{\mathcal{S}}_{t+1}(s_{0},\pi)=\{f(s,\pi(o))|$ $\|o-g(s,z)\|\leq\epsilon,$
@@
Proof.
Suppose $\mathcal{S}_{t}\subseteq\hat{\mathcal{S}}_{t}$ and let $s\in\mathcal{S}_{t}$ and $o=h(s,\omega)$ for some $\omega\in\Omega$. Then $s\in\hat{\mathcal{S}}_{t}(s_{0},\pi)$ and by Assumption (1), $\|o-g(s,z)\|\leq\epsilon$ for some $z\in\mathcal{Z}$. Consequently, $f(s,\pi(o))\in\hat{\mathcal{S}}_{t+1}(s_{0},\pi)$. Since $\hat{\mathcal{S}}_{0}(s_{0},\pi)=\mathcal{S}_{0}(s_{0},\pi)=\{s_{0}\}$, the result follows by induction. ∎
@@
Theorem 2.
Suppose that $\{s_{i}\}_{i=1}^{N}$ i.i.d. from $\mathcal{D}$ and let $V=\{s_{i}|P(\mathrm{Reach}_{K}(s_{i},\pi,g))\}$ be the subset of sampled initial states for which the reachable set is safe. Then under Assumption (1),
@@
with probability at least $1-\delta$.
Proof.
Let $\alpha=\Pr_{s\sim\mathcal{D}}\left[P(\mathrm{Reach}_{K}(s,\pi,g))\right]$ and $\hat{\alpha}=\frac{|V|}{N}$. By the Chernoff-Hoeffding bound, $\Pr\left(|\hat{\alpha}-\alpha|\geq\epsilon\right)\leq 2e^{-2N\epsilon^{2}}$, where the probability is with respect to datasets of $N$ initial states. Letting $\delta=2e^{-2N\epsilon^{2}}$, we obtain the confidence bound: $\Pr\left(\alpha\geq\hat{\alpha}-\sqrt{\frac{1}{2N}\log\frac{2}{\delta}}\right)\geq 1-\delta$. Finally, since by Theorem 1, $P(\mathrm{Reach}_{K}(s_{0},\pi,g))\Rightarrow P(\mathrm{Reach}_{K}(s_{0},\pi))$, the result follows. ∎
@@
4 Learning-Based Synthesis of Provably Safe Vision-Based Control
@@
The previous section shows that this still enables rigorous semi-probabilistic verification. Additionally, we consider a special case in which safety properties are tied to a scalar safety score (for example, cross-track error in lane following). In particular, let $\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$ be a safety score function over the reachable set, with $P(\mathcal{T})$ translating the safety score into a predicate (e.g., error exceeds a predefined threshold). We assume that we can obtain differentiable bounds on $\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$ (e.g., if we use $\alpha,\beta$-CROWN (zhang2018efficient; xu2021fast; xu2020automatic)).
4.1 The Learning Framework
@@
A central challenge in synthesizing a provably safe policy $\pi_{\theta}$ in our setting arises from the involvement of high-dimensional images generated by the generator $g$ (as a proxy for the true perception model $h$), which serve as inputs to the controller. Our overall approach is as follows. First, we begin with a pre-trained controller $\pi_{\hat{\theta}}$ that is empirically safe, which we also use as the anchor controller to avoid sacrificing too much empirical performance as we train for safety verification. Next, starting with $\hat{\theta}$, we train (or fine-tune) $\pi_{\theta}$ to minimize $\sum_{i\in S_{l}}\mathcal{L}(s_{i},\theta),$ where $S_{l}$ is a set of initial states $s_{0}$ used in training which evolves over training iterations $l$, and $\mathcal{L}(\theta)$ an appropriate loss function. The central algorithm design questions thus amount to 1) the choice of the loss function, and 2) the problem of selecting data $S_{l}$ to use for training in each iteration, so as to ultimately obtain a provably (rather than merely empirically) safe policy. We address these questions next.
4.2 Loss Function
@@
We propose a loss function that integrates both a performance-preserving loss and safety loss as follows:
$\mathcal{L}(s_{0},\theta)=\lambda_{1}\mathcal{L}_{\text{perf}}(s_{i},\theta)+\lambda_{2}\mathcal{L}_{\text{safety}}(s_{i},\theta).$ (4)
@@
The performance loss $\mathcal{L}_{\text{perf}}$ aims to preserve the empirical performance with respect to the pre-trained anchor controller $\pi_{\hat{\theta}}$. Depending on the training method, it can be (1) $\ell_{2}$ loss between predictions and ground truth for supervised training, or (2) RL objective (e.g. policy gradient) to encourage high expected returns for RL training.
@@
Turning next to the safety loss, recall that we assume that safety is quantified by a safety score function $\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$. One candidate would simply be to use this score as part of the loss function. However, this is impractical, as it is typically intractable to compute at scale and to the extent that it can be done, the tools for doing so are not differentiable. However, neural network verification techniques exist which compute differentiable sound upper and lower bounds on this quantity, and these therefore make natural candidates to use in constructing a loss function. More precisely, let $\underline{\sigma}(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))\leq\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))\leq\overline{\sigma}(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$ (i.e., $\underline{\sigma}(\cdot)$ is the lower and $\overline{\sigma}(\cdot)$ the upper bound on $\sigma(\cdot)$). To simplify notation, we let $\sigma^{(i)}_{K}=\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$, with $\overline{\sigma}^{(i)}_{K}$ and $\underline{\sigma}^{(i)}_{K}$ the corresponding upper and lower bounds. Then we define the safety loss as
@@
4.3 Adaptive Training Data
Our adaptive training procedure performs gradient updates by sampling batches from an adaptive collection of training data $S_{l}$ which consists of two disjoint and fixed-size components: the set of random initial states $S_{0}$, and $S_{A}$, maintained as a priority queue, containing initial states for which safety is a challenge to satisfy. Specifically, when training starts, $S_{A}$ is empty, and we gradually populate $S_{A}$ during the warmup period by adding the mm% most challenging datapoints (in the sense detailed below) from each training batch to it. To ensure $S_{0}$ and $S_{A}$ remain disjoint, whenever a datapoint is added to $S_{A}$, it is also deleted from $S_{0}$, and we generate another datapoint uniformly randomly to add to $S_{0}$.